A 4-m uniform bar of 10-kg mass is inclined 30° with the horizontal. If it is supported vertically at its ends find the force of reaction at each support.


1
Expert's answer
2021-05-27T14:19:21-0400

"mg\cos30°=R_{ay}+R_{by}"


"mg\sin30°=R_{ax}"



"(mgl/2)\cdot\cos30°-R_{by}\cdot l=0\to"


"R_{by}=(mg/2)\cdot\cos30°=(10\cdot9.8/2)\cdot\cos30°=42.44\ (N)" .



"R_{ax}=mg\sin30°=49\ (N)" .



"(mgl/2)\cdot\cos30°-R_{ay}\cdot l=0\to"


"R_{ay}=(mg/2)\cdot\cos30°=(10\cdot9.8/2)\cdot\cos30°=42.44\ (N)"



****


"R_a=\sqrt{42.44^2+49^2}=64.82\ (N)" . Answer


"R_b=42.44\ (N)" . Answer